OpenStudy (anonymous):

can someone show me how to use the product rule to find(x^2+1)(X^2-2x)

OpenStudy (anonymous):

sorry to diffferentation the function

OpenStudy (anonymous):

Recall the product rule first of all\[dy/dx=f \prime(x)g(x)+f(x)g \prime(x)\]Let \[f(x)=x^2+1\]\[g(x)=x^2-2x\]Then we get\[f \prime(x)=2x\]\[g \prime(x)=2x-2\]Substitute these into the product rule formula\[dy/dx=2x(x^2-2x)+(x^2+1)(2x-2)\]Expand and simplify if necessary

OpenStudy (anonymous):

well with the product rule you keep the first * the derivative of the second + keep the second * the derivative of the first.

OpenStudy (anonymous):

thanks think i got it now

OpenStudy (anonymous):

so you make them into derivivites as the first step all the time?

OpenStudy (anonymous):

Yep, look for 2 functions that are multiplied together and set one of them as f(x) and the other as g(x). Find their derivatives and use the formula. Should work every time.

OpenStudy (anonymous):

ooh got it now